# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ D.2 Proofs for robust characterization
$\overline{V}_{no}(\pi)$ $\coloneqq\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)q_{tr}(X)\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\right\}$
@@
$=\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[Y\mid T=t,X]q_{tr}(X)\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\right\}$
@@
$\beta_{t\mid r}(a)\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t,A=a],\beta_{t\mid r}\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t]$
@@
$={\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\{c_{t}\beta_{t\mid r}\}\pi_{r}(X)\mathbb{E}[Y\mid T=t,X]\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\colon\underline{B}\leq c_{1}\leq\overline{B},c_{0}=1-c_{1}\right\}$
$={\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\{c_{t}\beta_{t\mid r}\}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]\colon\underline{B}\leq c_{1}\leq\overline{B},c_{0}=1-c_{1}\right\}$ (LOTE)
$=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[c^{*}_{rt}\beta_{t\mid r}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}\left[X\in\mathcal{X}^{\text{nov}}\right]]]$
@@
Proof of proposition˜6.
$\min_{\pi}\;$ $\mathbb{E}[c(\pi,T(\pi),Y(\pi))\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]+\mathbb{E}[c(\pi,T(\pi),Y(\pi))\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]$ (7)
$\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=b]$ (8)
$\qquad\;\;+\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=b]\leq\epsilon,\forall q_{r1}\in\mathcal{U}$ (9)
@@
$t-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{r1}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}],\forall q_{r1}\in\mathcal{U}$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)q_{r1}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)q_{r1}(x,b))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,\forall q_{r1}\in\mathcal{U}$
@@
$t-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{r1}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}],\forall q_{r1}\in\mathcal{U_{1}}$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)q_{r1}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)q_{r1}(x,b))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,\forall q_{r1}\in\mathcal{U_{2}}$
@@
$t-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{r1}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}],\forall q_{r1}\in\mathcal{U_{1}}$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
We define
$\delta_{r1}(x,u)=\frac{2(q_{r1}(x,u)-\underline{B}_{r}(x,u))}{\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)}-(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)),$
then
$\{\underline{B}_{r}(x,u)\leq q_{r1}(x,u)\leq\overline{B}_{r}(x,u)\}\implies\{\lVert\delta_{r1}(x,u)\rVert_{\infty}\leq 1\}$
and
$q_{r1}(x,u)=\underline{B}_{r}(x,u)+\frac{1}{2}(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u))(\delta_{r1}(x,u)+1).$
For brevity we denote $\Delta B=(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u))$, so
@@
$t+\min_{\begin{subarray}{c}\lVert\delta_{r1}(x,u)\rVert_{\infty}\leq 1\\ r\in\{0,1\},u\in\{a,b\}\end{subarray}}\left\{-\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}\frac{1}{2}\Delta B(x,u)\delta_{r1}(x,u)dx\right\}$
$\qquad\qquad-c_{1}(\pi)\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}]$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,$
where
$c_{1}(\pi)=\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}(\underline{B}_{r}(x,u)+\frac{1}{2}(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)))dx$
@@
$t+\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}-\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\frac{1}{2}\Delta B(x,u)dx-c_{1}(\pi)\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}]$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
Undoing the epigraph transformation, we obtain:
$\min_{\pi}\;$ $V_{ov}(\pi)+\mathbb{E}[\mu_{0}]+c_{1}(\pi)+\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\frac{1}{2}\Delta B(x,u)dx$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
and simplifying the absolute value:
$\min_{\pi}\;$ $V_{ov}(\pi)+\mathbb{E}[\mu_{0}]+c_{1}(\pi)+\int_{x\in\mathcal{X}^{\text{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\frac{1}{2}\Delta B(x,u)dx$
$\int_{x\in\mathcal{X}^{\text{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
D.3 Proofs for general fairness-constrained policy optimization algorithm and analysis
@@
$g_{j}(O;Q)$ $=\mathbb{E}_{\pi\sim Q}[g_{j}(O;\pi)\mid O,\mathcal{E}_{j}]$
$h_{j}(Q)$ $=\mathbb{E}[g_{j}(O;Q)\mid\mathcal{E}_{j}]$
$\hat{h}^{1}_{j}(Q)$ $=\mathbb{E}_{n_{1}}[g_{j}(O;Q)\mid\mathcal{E}_{j}]$
D.3.1 Preliminaries: results from other works used without proof
Theorem 6 (Theorem 3, [Agarwal et al., 2018] (saddle point generalization bound for LABEL:alg-metaalg1) ).
Let $\xi:=\max_{h}\|{M}\hat{{\mu}}(h)-\hat{{c}}\|_{\infty}$. Suppose Assumption˜7 holds for $C^{\prime}\geq 2C+2+\sqrt{\ln(4/\delta)/2}$, where $\delta>0$. Let $Q^{\star}$ minimize $V(Q)$ subject to ${M}{h}(Q)\leq{c}$. Then Algorithm 1 with $\nu\propto n^{-\alpha},B\propto n^{\alpha}$ and $\omega\propto\xi^{-2}n^{-2\alpha}$ terminates in $O\left(\xi^{2}n^{4\alpha}\ln|\mathcal{K}|\right)$ iterations and returns $\hat{Q}$. If $np_{j}^{\star}\geq 8\log(2/\delta)$ for all $j$, then with probability at least $1-(|\mathcal{J}|+1)\delta$ then for all $k$, $\hat{Q}$ satisfies:
$V(\hat{Q})\leq V\left(Q^{\star}\right)+\widetilde{O}\left(n^{-\alpha}\right)$
$\gamma_{k}(\widehat{Q})\leq d_{k}+\frac{1+2\nu}{B}+\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\widetilde{O}\left(\left(np_{j}^{\star}\right)^{-\alpha}\right)$

### v5

@@ 12.3.1 Additional results
Let $\mathcal{U}_{\text{cbnd}}\coloneqq\left\{q_{t\mid r}(x^{\prime})\colon\underline{B}\leq q_{1\mid r}(x^{\prime})\leq\overline{B}\right\}$ and $\mathcal{U}=\mathcal{U}_{q_{t\mid r}}\cap\mathcal{U}_{\text{cbnd}}.$ Define $\bar{q}_{t\mid r}\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t].$If $T\in\{0,1\},$
$\overline{V}_{no}(\pi)$ $=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[c^{*}_{rt}\bar{q}_{t\mid r}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}_{\mathrm{nov}}(X)]],$
$\text{ where }\textstyle c^{*}_{rt}=\overline{B}\mathbb{I}\left[t=1\right]+\underline{B}\mathbb{I}\left[t=0\right]\text{ if }\mathbb{E}[Y\pi_{r}(X)\mid T=t]\geq 0,\text{ and }c^{*}_{rt}=0\text{ otherwise. }$
Proof 12.8
$V(\pi)=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[Y(t)\mathbb{I}\left[T(r)=t\right]\mid R=r,X]]$
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[Y(t)\mid R=r,X]P(T(r)=t\mid R=r,X)]$ unconf.
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[Y(t)\mid X]P(T(r)=t\mid R=r,X)]$ Section 3 (ER) (11)
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\mathbb{E}\left[Y(t)\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}\mid X\right]P(T(r)=t\mid R=r,X)\right]$ unconf.
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\mathbb{E}\left[Y(t)\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}+\left(1-\frac{T}{p_{t}(X)}\right)\mu_{t}(X)\mid X\right]p_{t\mid r}(X)\right\}\right]$ control variate
$=$ $\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}\left[\pi_{r}(X)\left\{\left\{Y(t)\frac{\mathbb{I}\left[T(r)=t\right]}{p_{t}(X)}+\left(1-\frac{T}{p_{t}(X)}\right)\mu_{t}(X)\right\}p_{t\mid r}(X)\right\}\right]$ (LOTE)
where $p_{t}(X)=P(T=t\mid X)$ (marginally over $R$ in the observational data) and (LOTE) is an abbreviation for the law of total expectation.
Proof 12.9
Proof of Lemma 12.7
@@
$\overline{V}_{no}(\pi)$ $\coloneqq\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)q_{tr}(X)\mathbb{I}_{\mathrm{nov}}(X)]]\right\}$
$=\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mathbb{E}[Y\mid T=t,X]q_{tr}(X)\mathbb{I}_{\mathrm{nov}}(X)]]\right\}$
@@
$\bar{q}_{t\mid r}(a)\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t,A=a],\bar{q}_{t\mid r}\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t]$
@@
$={\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\{c_{t}\bar{q}_{t\mid r}\}\pi_{r}(X)\mathbb{E}[Y\mid T=t,X]\mathbb{I}_{\mathrm{nov}}(X)]]\colon\underline{B}\leq c_{1}\leq\overline{B},c_{0}=1-c_{1}\right\}$
$={\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\{c_{t}\bar{q}_{t\mid r}\}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}_{\mathrm{nov}}(X)]]\colon\underline{B}\leq c_{1}\leq\overline{B},c_{0}=1-c_{1}\right\}$ (LOTE)
$=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[c^{*}_{rt}\bar{q}_{t\mid r}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}_{\mathrm{nov}}(X)]]$
@@
Proof 12.10
Proof of proposition 5.5
Remark 12.11
Remark 3. Proposition 5.5 in the main text is stated in value-maximization form. For the derivation below we work with the equivalent cost objective $C(\pi):=-\underline{V}_{\mathrm{rob}}(\pi)$ and convert back by multiplying the objective by 1-1.
$\min_{\pi}\;$ $C(\pi)$
s.t. $\mathbb{E}[T(\pi)\mathbb{I}_{\mathrm{ov}}(X)\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}_{\mathrm{ov}}(X)\mid A=b]$
$\qquad+\mathbb{E}[T(\pi)\mathbb{I}_{\mathrm{nov}}(X)\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}_{\mathrm{nov}}(X)\mid A=b]\leq\epsilon,\quad\forall q_{1\mid r}\in\mathcal{U}$ (12)
@@
$t-\int_{x\in\mathcal{X}_{\mathrm{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{1\mid r}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)],\forall q_{1\mid r}\in\mathcal{U}$
$\int_{x\in\mathcal{X}_{\mathrm{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)q_{1\mid r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)q_{1\mid r}(x,b))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,\forall q_{1\mid r}\in\mathcal{U}$
@@
$t-\int_{x\in\mathcal{X}_{\mathrm{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{1\mid r}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)],\forall q_{1\mid r}\in\mathcal{U_{1}}$
$\int_{x\in\mathcal{X}_{\mathrm{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)q_{1\mid r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)q_{1\mid r}(x,b))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,\forall q_{1\mid r}\in\mathcal{U_{2}}$
@@
$t-\int_{x\in\mathcal{X}_{\mathrm{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}q_{1\mid r}(x,u)\}dx\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)],\forall q_{1\mid r}\in\mathcal{U_{1}}$
$\int_{x\in\mathcal{X}_{\mathrm{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
Define midpoint and radius $m_{r}(x,u):=\tfrac{1}{2}(\underline{B}_{r}(x,u)+\overline{B}_{r}(x,u))$ and $\rho_{r}(x,u):=\tfrac{1}{2}(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u))$, and reparameterize
$q_{1\mid r}(x,u)=m_{r}(x,u)+\rho_{r}(x,u)\,\varphi_{1\mid r}(x,u),\qquad\|\varphi_{1\mid r}(x,u)\|_{\infty}\leq 1,$
so that $\underline{B}_{r}\leq q_{1\mid r}\leq\overline{B}_{r}\iff\|\varphi_{1\mid r}\|_{\infty}\leq 1$. Substituting,
@@
$t+\min_{\begin{subarray}{c}\lVert\varphi_{1\mid r}(x,u)\rVert_{\infty}\leq 1\\ r\in\{0,1\},u\in\{a,b\}\end{subarray}}\left\{-\int_{x\in\mathcal{X}_{\mathrm{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}\{g_{r}(x,u)\pi_{r}(x,u)f(x,u)\}\rho_{r}(x,u)\varphi_{1\mid r}(x,u)dx\right\}$
$\qquad\qquad-\sum_{r}\mathbb{E}\!\left[\pi_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\,m_{r}(X,A)\tau(X,A)\right]\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)]$
$\int_{x\in\mathcal{X}_{\mathrm{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon,$
@@
$t+\int_{x\in\mathcal{X}_{\mathrm{nov}}}\sum_{u\in\{a,b\}}\sum_{r\in\{0,1\}}-\left|g_{r}(x,u)\pi_{r}(x,u)f(x,u)\right|\rho_{r}(x,u)dx-\sum_{r}\mathbb{E}\!\left[\pi_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\,m_{r}(X,A)\tau(X,A)\right]\geq V_{ov}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)]$
$\int_{x\in\mathcal{X}_{\mathrm{nov}}}\{f(x\mid a)(\sum_{r}\pi_{r}(x,a)\overline{B}_{r}(x,a))-f(x\mid b)(\sum_{r}\pi_{r}(x,b)\underline{B}_{r}(x,u))\}+\mathbb{E}[\Delta_{ov}T(\pi)]\leq\epsilon$
Multiplying the objective by 1-1 converts this equivalent cost-minimization program back into the value-maximization statement of Proposition 5.5 in the main text. Equivalently, and matching the main-text statement, we obtain:
$\max_{\pi}\;$ $V_{\mathrm{ov}}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)]+c_{1}(\pi)-\sum_{r\in\{0,1\}}\mathbb{E}\!\left[|\tau(X,A)|\,\pi_{r}(X,A)\,\rho_{r}(X,A)\,\mathbb{I}_{\mathrm{nov}}(X)\right]$
s.t. $\sum_{r\in\{0,1\}}\Big\{\mathbb{E}[\pi_{r}(X,A)\overline{B}_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\mid A=a]-\mathbb{E}[\pi_{r}(X,A)\underline{B}_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\mid A=b]\Big\}+\Delta_{\mathrm{ov}}^{T}(\pi)\leq\epsilon.$
12.4 Proofs for general fairness-constrained policy optimization algorithm and analysis
@@
$\delta_{k}(O;Q)$ $=\mathbb{E}_{\pi\sim Q}[\delta_{k}(O;\pi)\mid O]$
$\Delta_{k}(Q)$ $=\mathbb{E}[\delta_{k}(O;Q)]$
$\hat{\Delta}^{1}_{k}(Q)$ $=\mathbb{E}_{n_{1}}[\delta_{k}(O;Q)]$
12.4.1 Preliminaries: results from other works used without proof
The original reductions theorem is stated for cost minimization. We apply it to the equivalent cost objective $\tilde{V}=-V$ and state the resulting sign-flipped corollary below in our value-maximization notation.
Theorem 12.12 (Theorem 3, [Agarwal et al., 2018] (saddle point generalization bound for LABEL:alg-metaalg1) )
Let $\xi:=\sup_{Q\in\mathcal{Q}}\|\hat{\Delta}(Q)-\Delta(Q)\|_{\infty}$. Suppose Section 5.2 holds for $C^{\prime}\geq 2C+2+\sqrt{\ln(4/\delta)/2}$, where $\delta>0$. Let $Q^{\star}$ maximize $V(Q)$ subject to $\Delta(Q)\leq d$. Then Algorithm 1 with $\nu\propto n^{-\alpha},B\propto n^{\alpha}$ and $\omega\propto\xi^{-2}n^{-2\alpha}$ terminates in $O\left(\xi^{2}n^{4\alpha}\ln|\mathcal{K}|\right)$ iterations and returns $\hat{Q}$. If $np_{j}^{\star}\geq 8\log(2/\delta)$ for all $j$, then with probability at least $1-(|\mathcal{J}|+1)\delta$ then for all $k$, $\hat{Q}$ satisfies:
$V(\hat{Q})\geq V\left(Q^{\star}\right)-\widetilde{O}\left(n^{-\alpha}\right)$
$\Delta_{k}(\widehat{Q})\leq d_{k}+\frac{1+2\nu}{B}+\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\widetilde{O}\left(\left(np_{j}^{\star}\right)^{-\alpha}\right)$